OpenStudy (anonymous):

Use the elimination method to solve the system for x and y. 2x+3y=19 8x-3y=-14 Select one: a. (1/2, 6) b. (6, 1/2) c. (3.3, 4) d. (0.5, -6)

OpenStudy (anonymous):

2x+3y=19 + 8x-3y=-14 _______________ 10x=5 x=5/10 x=1/2 substitute back into 1: 2(1/2)+3y=19 1+3y=19 3y=18 y=6 Therefore answer is a. (1/2,6)

OpenStudy (anonymous):

:)thanks(: